OpenStudy (anonymous):

Find missing numbers 1/4 c = ?tbsp 8hr=?sec

OpenStudy (jlastino):

Assuming c is cups 1/4 cups =4 tablespoons 8hr = 28 800

OpenStudy (anonymous):

How do you get the answer?

OpenStudy (jlastino):

well 1cup is equal to 16 table spoons so we 1/4 cup would be equal to 16/4 table spoons 1 hour = 60 minutes 1 minutes = 60 seconds 1 hour =3600 seconds multiply by this by 8 3600*8=28 800